# example.bib

### v1

@@ II-A Learning from Mixed-Quality Demonstrations
Several studies address the challenge of learning from mixed-quality demonstrations by introducing confidence weights to assess the optimality of state-action pairs [ileed, demodice, bcnd, swbt, dwbc, cail]. BCND [bcnd] uses the policy from previous steps as weights for iterative refinement. DWBC [dwbc] and ILEED [ileed] enhance policy performance through joint training with additional components: a discriminator and a state-encoder, respectively. However, these methods risk guiding the model towards incorrect distributions or increasing sensitivity to hyperparameters and training instability. Concurrently, a recent approach [yin2024offline] constructs a state graph from the mixed-quality demonstrations, linking nearby states in visual representation space and consecutive states in demonstrations to determine confidence through graph search. Still, the effectiveness depends on tuning the tolerance range for nearby states for optimal performance.
@@
The demonstration selection method decouples data processing from policy learning, simplifying the handling of mixed-quality data and enabling policy training without modifications. L2D [l2d] employs preference learning on augmented trajectories to filter out low-quality demonstrations, while ELICIT [elicit] aligns demonstrations with the expert policy but may neglect the usefulness of low-quality data. PUBC [pubc] creates negative samples by mismatching state-action pairs and updates the positive dataset iteratively through adaptive voting, but it overlooks the fact that state-action pairs hold more significance within the context of the entire sequence. AWE [awe] applies dynamic programming to smooth expert trajectories, potentially preserving outliers in mixed-quality data. These methods reflect the difficulty of utilizing mixed-quality demonstrations without oversimplifying the data. Moreover, simply discarding the low-quality data [elicit, l2d, pubc] can lead to inadequate utilization of the mixed-quality data and limit the downstream policy performance.
@@ III-A Problem Formulation
Consider an offline dataset $\mathcal{D}=\{\tau_{1},\tau_{2},\ldots,\tau_{N}\}$ containing $N$ mixed-quality demonstrations for a robotic manipulation task. Each demonstration $\tau$ is a trajectory $\left(o_{1},a_{1},o_{2},a_{2},\ldots,o_{T},a_{T}\right)$, where $t\in{1,2,\ldots,T}$ represents the timestep, $o_{t}$ is the observation, and $a_{t}$ is the action taken by the robot. Additionally, a small offline dataset $\mathcal{D}^{*}=\{\tau^{*}_{1},\tau^{*}_{2},\cdots,\tau^{*}_{M}\}$ of $M\ll N$ expert demonstrations is also provided for reference and cannot be used for policy learning. In some cases, each demonstration $\tau_{i}$ in $\mathcal{D}$ may have a subjective quality label $l_{i}$, which is assumed to be inaccessible in our problem setting.
@@ III-D Trajectory Optimization
Since the demonstration segment $\tau$ contains semantically consistent robot actions, we propose a greedy algorithm (Alg. 1) to approximate the optimal path $\tau^{\prime}$ by eliminating points with excessive spatial deviation. For $\tau\in\mathcal{D}_{s,-}$, we start with $\tau^{\prime}=\{(o_{1},a_{1})\}$ and iteratively select the next waypoint. Define $\mathbf{e}_{i,j}:=e_{i}-e_{j}$ as the vector from $e_{j}$ to $e_{i}$. We construct a candidate set $K$ for the next waypoint $(o_{k},a_{k})$ using a threshold $\delta_{\theta}$ for the maximum angle between $\mathbf{e}_{k,j}$ and $\mathbf{e}_{T,j}$, where $T$ is the ending timestep of the segment. This process filters out unnecessary actions that deviate significantly from the action trend within the segment. We then choose $(o_{k},a_{k})$ with the minimum distance $\left|\mathbf{e}_{k,j}\right|$ as the next waypoint and add it to $\tau^{\prime}$. The selection process continues until $(o_{T},a_{T})$ is included, forming the approximated optimal path $\tau^{\prime}$ for the low-quality segment $\tau$.
@@ IV-B Simulation Results
S2I can improve the performance of various downstream policies by optimizing the mixed-quality demonstration data (Q1). Fig. 6, Tab. I and Tab. VIII illustrate that S2I consistently outperforms baseline methods in simulation environments, improving the performance of various robotic manipulation policies on mixed-quality demonstrations. This underscores the effectiveness of S2I in handling such data. In line with [dp], we also report the average performance of the last 10 checkpoints to provide a representative assessment of policy performance during training and mitigate the influence of outliers. S2I continues to surpass previous methods by a considerable margin ($\color[rgb]{0.76,0.23,0.13}\definecolor[named]{pgfstrokecolor}{rgb}{0.76,0.23,0.13}\sim+20\%$), demonstrating its ability to filter and optimize low-quality segments, resulting in substantial performance gains.
@@ IV-C Ablations
Segment-level selection and trajectory optimization (with action relabeling) are both crucial for performance improvement (Q3). We ablate several design choices, including the selection level and the trajectory optimization module. For demonstration-level selection, we assess demonstration quality by averaging the quality results of each segment instead of using the operator expertise labels provided in the dataset. The ablation results are shown in Tab. III. If we apply no selection, performance drops when only trajectory optimization is applied ($\color[rgb]{0.01,0.75,0.24}\definecolor[named]{pgfstrokecolor}{rgb}{0.01,0.75,0.24}-27.48\%$), as high-quality segments are relatively optimal and thus do not require further optimization. By selecting at the demonstration or segment level, though data efficiency becomes low, the policy performance can be improved marginally ($\color[rgb]{0.76,0.23,0.13}\definecolor[named]{pgfstrokecolor}{rgb}{0.76,0.23,0.13}+0.81\%$ and $\color[rgb]{0.76,0.23,0.13}\definecolor[named]{pgfstrokecolor}{rgb}{0.76,0.23,0.13}+1.84\%$). Introducing trajectory optimization with action relabeling boosts both dataset utilization and performance. Applying trajectory optimization at the segment level achieves the best results ($\color[rgb]{0.76,0.23,0.13}\definecolor[named]{pgfstrokecolor}{rgb}{0.76,0.23,0.13}+17.07\%$), maximizing data usage and enhancing model performance by refining the actions in low-quality data. The results validate the design choices of S2I.
@@ IV-D Real-World Results
S2I effectively optimizes real-world mixed-quality demonstration data, enhancing the performance of various downstream policies in real-world experiments (Q4). The results shown in Tab IV reveal that S2I can significantly improve performance for various downstream policies across different tasks, especially for those with initially low success rates. Hence, S2I is also effective in utilizing real-world mixed-quality demonstrations and can contribute to a better robotic manipulation policy.
@@ V Conclusion
This paper introduces “Select Segments to Imitate” (S2I), a framework designed to effectively use mixed-quality demonstrations for robotic manipulation policy learning by focusing on segment-level selection and optimization. S2I extracts high-quality segments via contrastive learning and refines lower-quality ones through trajectory optimization and action relabeling, ensuring effective learning of corrective behaviors while fully utilizing the demonstration dataset. Extensive experiments in simulated and real environments show the ability of S2I to optimize demonstrations and boost robotic manipulation policy performance. While S2I shows promising results across various tasks and policies, it may struggle in handling robot trajectories with complex rotations. Future work includes addressing this issue, applying S2I to large-scale robotic datasets [rh20t], and extending S2I to optimize language-guided demonstrations.
@@ Acknowledgement
We would like to thank Chenxi Wang for his help during the data collection process.

### v2

@@ II-A Learning from Mixed-Quality Demonstrations
Several studies address the challenge of learning from mixed-quality demonstrations by introducing confidence weights to assess the optimality of state-action pairs [ileed, demodice, bcnd, swbt, wu2019imitation, dwbc, cail]. BCND [bcnd] uses the policy from previous steps as weights for iterative refinement. DWBC [dwbc] and ILEED [ileed] enhance policy performance through joint training with additional components: a discriminator and a state-encoder, respectively. However, these methods risk guiding the model towards incorrect distributions or increasing sensitivity to hyperparameters and training instability. Concurrently, a recent approach [yin2024offline] constructs a state graph from the mixed-quality demonstrations, linking nearby states in visual representation space and consecutive states in demonstrations to determine confidence through graph search. Still, the effectiveness depends on carefully tuning the tolerance range for nearby states.
@@
The demonstration selection method decouples data processing from policy learning, simplifying the handling of mixed-quality data. L2D [l2d] employs preference learning to filter out low-quality demonstrations, while ELICIT [elicit] aligns demonstrations with expert policy but may neglect the usefulness of low-quality data. PUBC [pubc] creates negative samples by mismatching state-action pairs and updates the positive dataset iteratively through adaptive voting, , overlooking the significance of sequences. AWE [awe] applies dynamic programming to smooth expert trajectories, potentially preserving outliers in mixed-quality data. These methods reflect the difficulty of utilizing mixed-quality demonstrations without oversimplifying the data. Moreover, simply discarding the low-quality data [bu2024aligning, elicit, l2d, pubc] can lead to inadequate data utilization and limit downstream policy performance.
@@ III-A Problem Formulation
Consider an offline dataset $\mathcal{D}=\{\tau_{1},\tau_{2},\ldots,\tau_{N}\}$ containing $N$ mixed-quality demonstrations for a robotic manipulation task. Each demonstration $\tau$ is a trajectory $\left(o_{1},a_{1},o_{2},a_{2},\ldots,o_{T},a_{T}\right)$, where $t\in\{1,2,\ldots,T\}$ represents the timestep, $o_{t}$ is the observation, and $a_{t}$ is the action taken by the robot. Additionally, a small offline dataset $\mathcal{D}^{*}=\{\tau^{*}_{1},\tau^{*}_{2},\cdots,\tau^{*}_{M}\}$ of $M\ll N$ expert demonstrations is also provided for reference and cannot be used for policy learning. In some cases, each demonstration $\tau_{i}$ in $\mathcal{D}$ may have a subjective quality label $l_{i}$, which is assumed to be inaccessible in our problem setting.
@@ III-D Trajectory Optimization
Since the demonstration segment $\tau$ contains semantically consistent robot actions, we propose a greedy algorithm (Alg. 1) to approximate the optimal path $\tau^{\prime}$ by eliminating points with excessive spatial deviation. For $\tau\in\mathcal{D}_{s,-}$, we start with $\tau^{\prime}=\{(o_{1},a_{1})\}$ and iteratively select the next waypoint. Define $\mathbf{e}_{i,j}:=e_{i}-e_{j}$ as the vector from $e_{j}$ to $e_{i}$, where $e_{i}$ represents the end-effector position at point $i$. We construct a candidate set $K$ for the next waypoint $(o_{k},a_{k})$ using a threshold $\delta_{\theta}$ for the maximum angle between $\mathbf{e}_{k,j}$ and $\mathbf{e}_{T,j}$, where $T$ is the ending timestep of the segment. This process filters out unnecessary actions that deviate significantly from the action trend within the segment. We then choose $(o_{k},a_{k})$ with the minimum distance $\left|\mathbf{e}_{k,j}\right|$ as the next waypoint and add it to $\tau^{\prime}$. The selection process continues until $(o_{T},a_{T})$ is included, forming the approximated optimal path $\tau^{\prime}$ for the low-quality segment $\tau$.
@@ IV-B Simulation Results
S2I can improve the performance of various downstream policies by optimizing the mixed-quality demonstration data (Q1). Fig. 6, Tab. I and Tab. VIII illustrate that S2I consistently outperforms baseline methods in simulation environments. This underscores the effectiveness of S2I in handling such data. In line with [dp], we also report the average performance of the last 10 checkpoints to provide a representative assessment during training. S2I continues to surpass previous methods by a considerable margin ($\color[rgb]{0.76,0.23,0.13}\definecolor[named]{pgfstrokecolor}{rgb}{0.76,0.23,0.13}\sim+20\%$), demonstrating its ability to filter and optimize low-quality segments for substantial performance gains.
@@ IV-C Ablations
Segment-level selection and trajectory optimization (with action relabeling) are both crucial for performance improvement (Q3). We ablate the selection level and the trajectory optimization module. For demonstration-level selection, we assess demonstration quality by averaging the quality results of each segment. The ablation results in Tab. III show that without selection, performance drops when only trajectory optimization is applied ($\color[rgb]{0.01,0.75,0.24}\definecolor[named]{pgfstrokecolor}{rgb}{0.01,0.75,0.24}-27.48\%$), as high-quality segments are already optimal. Selecting at demonstration or segment level improves performance marginally ($\color[rgb]{0.76,0.23,0.13}\definecolor[named]{pgfstrokecolor}{rgb}{0.76,0.23,0.13}+0.81\%$ and $\color[rgb]{0.76,0.23,0.13}\definecolor[named]{pgfstrokecolor}{rgb}{0.76,0.23,0.13}+1.84\%$). Introducing trajectory optimization with action relabeling boosts both dataset utilization and performance, with segment-level application achieving the best results ($\color[rgb]{0.76,0.23,0.13}\definecolor[named]{pgfstrokecolor}{rgb}{0.76,0.23,0.13}+17.07\%$) by refining actions in low-quality data. These results validate S2I’s design choices.
@@ IV-D Real-World Results
S2I effectively optimizes real-world mixed-quality demonstration data, enhancing the performance of various downstream policies in real-world experiments (Q4). The comprehensive results shown in Tab IV reveal that S2I can significantly improve performance for various downstream policies across different tasks, especially for those with initially low success rates. Hence, S2I is also consistently effective in utilizing real-world mixed-quality demonstrations and can substantially contribute to a better robotic manipulation policy.
@@ V Conclusion
This paper introduces “Select Segments to Imitate” (S2I), a framework designed to effectively use mixed-quality demonstrations for robotic manipulation policy learning by focusing on segment-level selection and optimization. S2I extracts high-quality segments via contrastive learning and refines lower-quality ones through trajectory optimization and action relabeling, ensuring effective learning of corrective behaviors while fully utilizing the demonstration dataset. Extensive experiments in both simulated and real-world environments demonstrate the robust ability of S2I to optimize demonstrations and significantly boost robotic manipulation policy performance. While S2I shows promising results across various tasks and policies, it may struggle in handling robot trajectories with complex rotations. Future work includes addressing this issue, applying S2I to large-scale robotic datasets [rh20t], and extending S2I to optimize language-guided demonstrations.
@@ Acknowledgement
This work was supported by the National Key Research and Development Project of China (No. 2022ZD0160102), the National Key Research and Development Project of China (No. 2021ZD0110704), Shanghai Artificial Intelligence Laboratory, XPLORER PRIZE grants.